# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

### v2

@@ B.3 Training Details
B.4 Formal Definitions of Metrics
@@
B.5 Policy Optimization Algorithm: GRPO
The policy is optimized using Group Relative Policy Optimization (GRPO) (grpo2024), a policy gradient method that operates without a learned value function. In GRPO, the advantage for a given response is computed by normalizing its reward against the statistics of a group of candidate responses sampled for the same prompt.
@@ B.6 Baseline Details
  • Offline Baseline: Tasks are pre-sorted once from easy to hard based on the success rates of external models (Qwen2.5-7B-Instruct, with Qwen3-8B for tie-breaking). This baseline represents a static curriculum and is used to benchmark our adaptive method against a fixed task sequence.
@@
  • Setting from MoPPS (qu2025can): This ablation uses $\lambda=0.0,\rho=0.0$ with posterior sampling, relying solely on explicit evidence for task selection. Under this configuration, our framework reduces exactly to the setting studied in qu2025can, enabling a direct evaluation of a purely explicit-evidence-based strategy. Note that qu2025can did not study the code or logic domains, nor did they provide principles for determining $\lambda$ ($1-\lambda$ in BOTS). We therefore reuse their hyperparameter choice for math ($\lambda=1$, i.e., $\lambda=0$ in BOTS) when evaluating other domains. This comparison is fair since no hyperparameter tuning is performed for either MoPPS or BOTS.
@@
  • Proxy for DOTS (sun2025improving): This ablation uses $\lambda=1.0,\rho=1.0$ with posterior sampling disabled, so task selection is driven entirely by implicit evidence from our estimator. It serves as a proxy for the approach in sun2025improving, though our construction of pseudo-counts differs, see Appendix E.2 for more details. This baseline allows us to evaluate the long-term efficacy of a strategy that relies almost exclusively on implicit evidence without direct corrective feedback. We note that sun2025improving additionally explored reusing historical trajectories, which lies beyond the scope of this paper. To ensure clarity, we isolate such tricks from DOTS and focus solely on the online task selection component.
Appendix C More Related Works
Reinforcement Finetuning (RFT).
Reinforcement finetuning (RFT) has become a pivotal paradigm for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning capabilities. Early approaches, commonly known as Reinforcement Learning from Human Feedback (RLHF) (ouyang2022rlhf), train a reward model to approximate human preferences and then optimize the policy using reinforcement learning algorithms such as PPO (2017ppo). Subsequent research introduced direct preference optimization methods such as DPO (2023dpo) and its variants KTO (2024kto), ORPO (2024orpo), and SimPO (2024simpo), which bypass the need to train an explicit reward model. More recently, researchers have observed that when rewards are automatically verifiable—such as in mathematics or code generation—RFT can substantially enhance reasoning performance, leading to the emergence of verifiable-RL frameworks like RLVR, where GRPO (grpo2024) serves as a PPO-style adaptation designed for such settings. Building on these foundations, variants like DAPO (yu2025dapo) and GSPO (2025gspo) further refine the optimization dynamics, while other studies explore entropy regularization (2025primeentropy; 2025entropy28) and reward shaping (2025seedgrpo; 2025trinity). Despite these algorithmic advances, a critical bottleneck remains—the data curriculum, i.e., how tasks are selected and presented during RFT, which crucially determines training efficiency.
@@ Appendix C More Related Works
The importance of data difficulty in RFT is increasingly recognized, leading to the creation of highly challenging benchmarks (2025bigmath; he2025deepmath; gao2025omni). Beyond dataset creation, data selection and curriculum design has also become a central topic of investigation. For Supervised Finetuning (SFT), a variety of data selection and curriculum strategies have been widely studied based on quality (2024deita; 2024superfilter), diversity (2025daar; 2024instag), or pre-assessed difficulty (2025mindgym; 2025limo). However, such static methods are ill-suited for the dynamic nature of RFT. This has motivated a focus on dynamic curricula that adapt to the model’s evolving capabilities, typically by leveraging the notion of task difficulty (cheng2025revisiting; 2025hardexamples; li2025can).
@@
Existing task selection strategies for RFT vary in their approach to leveraging task difficulty. An early line of work, directly inspired by curriculum learning, employs offline curricula that schedule tasks along a fixed easy-to-hard trajectory (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple). While simple and intuitive, these methods are non-adaptive and cannot respond to the model’s real-time learning progress. To address this, online selection strategies have emerged. A straightforward approach is sampling-based task filtering, which uses extra rollouts to evaluate and discard tasks that are too easy or too hard, thereby incurring significant computational overhead (yu2025dapo; bae2025online). To avoid this cost, recent works attempt to predict task success rates. Some frame task selection as a non-stationary multi-armed bandit (MAB) problem, but often treat tasks as independent arms, thus overlooking cross-task relationships (chen2025self; qu2025can). Others use a small reference set to predict the difficulty of other tasks via similarity kernels, but this still requires extra rollouts and discards valuable historical information (sun2025improving; wang2025dump).
@@
Appendix D Proofs
D.1 Proof for Proposition 3.2
[…]
Proof.
D.2 Proof for Proposition 3.4
[…]
Proof.
We drop the task index for clarity. From Equation (1) and Equation (3),
Appendix E Extended Discussion
E.1 Impact of $\lambda,\rho$ on Task Selection
Proposition 3.4 characterizes how the hyperparameters $\lambda$ and $\rho$ jointly determine the effective sample size of each task’s Beta posterior. Intuitively, the effective sample size $n_{t}=\alpha_{t}+\beta_{t}$ controls the confidence of the estimated probability of success, and hence governs the tradeoff between exploration and exploitation in task selection. When a task is repeatedly selected and directly evaluated, its posterior accumulates evidence and $n_{t}$ approaches the upper bound $n_{0}+\tfrac{1}{\lambda}n$. Conversely, when a task is rarely selected and only updated by implicit evidence, $n_{t}$ remains close to the lower bound $n_{0}+\tfrac{\rho}{\lambda}n$.
The role of $\lambda$ is to regulate the overall scale of $n_{t}$. A smaller $\lambda$ accelerates the discounting of historical information, leading to smaller effective sample sizes, higher posterior uncertainty, and thus more exploratory behavior in Thompson Sampling. In contrast, a larger $\lambda$ stabilizes the posterior by retaining more historical counts, biasing the strategy toward exploitation of tasks already known to be near the target probability of success.
The parameter $\rho$ specifically controls the lower bound of $n_{t}$. A smaller $\rho$ increases the uncertainty of tasks without direct evaluations, making it more likely for such tasks to be sampled for exploration even when their estimated probability of success is not close to the target. A larger $\rho$, by assigning more pseudo-counts, suppresses this uncertainty and reduces the chance of revisiting unevaluated tasks, favoring exploitation.
Together, $\lambda$ and $\rho$ provide fine-grained control over the exploration–exploitation balance: $\lambda$ sets the overall confidence level of the posterior across all tasks, while $\rho$ tunes the relative confidence between evaluated and unevaluated tasks. This interaction ensures that the framework adaptively selects tasks of appropriate difficulty as the model evolves, maintaining both robustness and efficiency in online task selection.
E.2 Attention-Based Adaptive Difficulty Estimate
@@
For example, sun2025improving proposed to use an attention-like kernel structure:
@@
E.3 Interpolation-Based Implicit Evidence
@@
Unlike kernel-based estimators (sun2025improving) (see Appendix E.2 for a brief introduction), which are confined to the convex hull of observed tasks, our interpolation-based estimator naturally supports extrapolation. For example, if the reference pair consists of a strong and a stronger model, the capability coefficient $\mu_{t}(\mathcal{B}_{t})$ defined in Section 3.3 may fall outside $[0,1]$, yielding extrapolated predictions. The clipping step in Equation (5) ensures these estimates remain within the feasible range $[0,1]$.
Choice of Reference Models.
The effectiveness of the implicit evidence mechanism depends critically on the choice of weak and strong reference models, which should ideally bracket the expected capability trajectory of the model under training. If the training model is substantially stronger than the strong reference or weaker than the weak reference, our method must rely on extrapolation, whose accuracy may degrade. In practice, we find that standard publicly available checkpoints—Qwen2.5-7B-Instruct (Yang2024Qwen25TR) and Qwen3-32B (yang2025qwen3)—provide a robust and practical basis, even when training significantly smaller models such as Qwen2.5-1.5B-Instruct (Yang2024Qwen25TR). A systematic ablation on the choice of reference models remains an important direction for future work.
@@
The simplicity of our interpolation-based approach, while a key strength, also entails two potential limitations. (i) Expressive power: The linear interpolation assumes a linear relationship between a model’s global capability and its per-task success rate. While our empirical results suggest this is a powerful approximation (see Appendix F.2 for a validation), the true learning dynamics of LLMs may be more complex. (ii) Distributional shift in capability estimation: The capability coefficient $\mu_{t}$ is estimated on the selected batch $\mathcal{B}_{t}$, which is not a uniform sample from the task pool. This introduces a bias, as the model is likely to perform better on this adaptively chosen batch than on the entire dataset. Consequently, $\mu_{t}$ might be an overestimate of the model’s true global capability. However, we argue this bias is not fatal: the primary role of $\mu_{t}$ is to track the progression of the model’s capability, and even a biased estimate can provide a valuable monotonic signal for this purpose, see Section 4.4 for empirical validation.
E.4 Discussion on Future Works
E.4.1 Generalization to Other Reward Distributions
@@
E.4.2 Self-Adaptive Update Rules
@@
E.4.3 Alternative Plug-In for Implicit Evidence
Our interpolation-based estimator provides an extremely lightweight way to generate implicit evidence without additional rollouts, but it is not the only possible choice. More expressive alternatives could be explored, such as kernel-based predictors (sun2025improving), task-embedding regressors, or small auxiliary models trained jointly with the main model. These alternatives may improve predictive accuracy, especially when the reference models poorly bracket the training model’s capability. However, they also introduce a trade-off: richer implicit evidence often requires higher computational and storage costs. Systematically characterizing this trade-off—between accuracy and efficiency in implicit evidence—remains an open research question and a promising direction for future work.
Appendix F Additional Experimental Results
F.1 Computational Overhead

### v3

@@ D.3 Trainining Details
D.4 Formal Definitions of Metrics
@@
D.5 Policy Optimization Algorithm: GRPO
The policy is optimized using Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a policy gradient method that operates without a learned value function. In GRPO, the advantage for a given response is computed by normalizing its reward against the statistics of a group of candidate responses sampled for the same prompt.
@@ D.6 Baseline Details
  • Offline Baseline: Tasks are pre-sorted once from easy to hard based on the success rates of external models (Qwen2.5-7B-Instruct, with Qwen3-32B-A3B for tie-breaking). This baseline represents a static curriculum and is used to benchmark our adaptive method against a fixed task sequence.
@@
  • Setting from MoPPS (Qu et al., 2025): This ablation uses $\lambda=0.0,\rho=0.0$ with posterior sampling, relying solely on explicit evidence for task selection. Under this configuration, our framework reduces exactly to the setting studied in Qu et al. (2025), enabling a direct evaluation of a purely explicit-evidence-based strategy. Note that Qu et al. (2025) did not study the code or logic domains, nor did they provide principles for determining $\lambda$ ($1-\lambda$ in BOTS). We therefore reuse their hyperparameter choice for math ($\lambda=1$, i.e., $\lambda=0$ in BOTS) when evaluating other domains. This comparison is fair since no hyperparameter tuning is performed for either MoPPS or BOTS.
@@
  • Proxy for DOTS (Sun et al., 2025): This ablation uses $\lambda=1.0,\rho=1.0$ with posterior sampling disabled, so task selection is driven entirely by implicit evidence from our estimator. It serves as a proxy for the approach in Sun et al. (2025), though our construction of pseudo-counts differs, see Appendix F.2 for more details. This baseline allows us to evaluate the long-term efficacy of a strategy that relies almost exclusively on implicit evidence without direct corrective feedback. We note that Sun et al. (2025) additionally explored reusing historical trajectories, which lies beyond the scope of this paper. To ensure clarity, we isolate such tricks from DOTS and focus solely on the online task selection component.
Appendix E More Related Works
Reinforcement Finetuning for LLMs
Reinforcement Finetuning (RFT) has become a pivotal technique for aligning LLMs with human values and enhancing their capabilities in complex reasoning tasks (OpenAI, 2023; Guo et al., 2025; Zeng et al., 2025; He et al., 2025a). Initial RLHF methods using PPO (Schulman et al., 2017) have been complemented by preference optimization techniques like DPO (Rafailov et al., 2023) and its variants such as KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024). More recently, attention has shifted to rejection-sampling-based finetuning, popularized by frameworks like GPRO (Shao et al., 2024) for its effectiveness in verifiable reasoning. This has spurred variants like DAPO (Yu et al., 2025) and GSPO (Zheng et al., 2025), with further work focusing on algorithmic refinements (Cui et al., 2025; Wang et al., 2025b) and reward shaping (Chen et al., 2025a; Pan et al., 2025). Despite these algorithmic advances, the data curriculum, that how tasks are selected and presented, remains a critical bottleneck for RFT efficiency.
@@ Appendix E More Related Works
The importance of data difficulty in RFT is increasingly recognized, leading to the creation of highly challenging benchmarks (Albalak et al., 2025; He et al., 2025b; Gao et al., 2025). Beyond dataset creation, data selection and curriculum design has also become a central topic of investigation. For Supervised Finetuning (SFT), a variety of data selection and curriculum strategies have been widely studied based on quality (Liu et al., 2024; Li et al., 2024a), diversity (Ling et al., 2025; Lu et al., 2024), or pre-assessed difficulty (Xu et al., 2025; Ye et al., 2025). However, such static methods are ill-suited for the dynamic nature of RFT. This has motivated a focus on dynamic curricula that adapt to the model’s evolving capabilities, typically by leveraging the notion of task difficulty (Cheng et al., 2025; Pikus et al., 2025; Li et al., 2025b; Toloubidokhti et al., 2023; Wang et al., 2025a).
@@
Existing task selection strategies for RFT vary in their approach to leveraging task difficulty. An early line of work, directly inspired by curriculum learning, employs offline curricula that schedule tasks along a fixed easy-to-hard trajectory (Parashar et al., 2025; Shen et al., 2025; Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a). While simple and intuitive, these methods are non-adaptive and cannot respond to the model’s real-time learning progress. To address this, online selection strategies have emerged. A straightforward approach is sampling-based task filtering, which uses extra rollouts to evaluate and discard tasks that are too easy or too hard, thereby incurring significant computational overhead (Yu et al., 2025; Bae et al., 2025). To avoid this cost, recent works attempt to predict task success rates. Some frame task selection as a non-stationary multi-armed bandit (MAB) problem, but often treat tasks as independent arms, thus overlooking cross-task relationships (Chen et al., 2025b; Qu et al., 2025). Others use a small reference set to predict the difficulty of other tasks via similarity kernels, but this still requires extra rollouts and discards valuable historical information (Sun et al., 2025; Wang et al., 2025c).
@@
Appendix F Extended Discussion
F.1 A toy example illustrating the role of $\lambda$ in difficulty estimation
Figure 4: A toy example illustrating the role of $\lambda$ in difficulty estimation. Ground Truth simulates a continuously evolving success probability, while Observations add Gaussian noise to mimic estimation errors from explicit/implicit evidence. We plot the posterior means obtained by BOTS over 50 update steps for three settings of $\lambda\in\{0.1,0.2,0.5\}$. A larger $\lambda$ (e.g., 0.5) reacts quickly to rapid changes in the underlying difficulty but is highly sensitive to noise. A smaller $\lambda$ (e.g., 0.1) yields smoother and more stable estimates but lags behind when the true difficulty increases rapidly. The intermediate setting $\lambda=0.2$ balances the two behaviors, offering better stability than $\lambda=0.5$ and better adaptivity than $\lambda=0.1$.
F.2 Attention-Based Adaptive Difficulty Estimate
@@
For example, Sun et al. (2025) proposed to use an attention-like kernel structure:
@@
F.3 Interpolation-Based Implicit Evidence
When linear interpolation is exact and why it becomes imperfect in practice.
Assumption 1 (Linear Interpolation Assumption).
Let $p(\mathcal{T})$, $p_{w}^{\mathrm{ref}}(\mathcal{T})$, and $p_{s}^{\mathrm{ref}}(\mathcal{T})$ denote the success probability of the current model, the weak reference model, and the strong reference model, respectively. We assume that all tasks share a common interpolation coefficient, i.e., for any pair of tasks $\mathcal{T}^{1},\mathcal{T}^{2}$,
$\frac{p(\mathcal{T}^{1})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{1})}{p_{s}^{\mathrm{ref}}(\mathcal{T}^{1})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{1})}=\frac{p(\mathcal{T}^{2})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{2})}{p_{s}^{\mathrm{ref}}(\mathcal{T}^{2})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{2})}.$
We denote the (task-invariant) interpolation coefficient by $\mu^{\star}$.
This assumption states that the current model’s success probabilities lie on the same affine transformation of the weak and strong reference models; hence linear interpolation is a perfect estimator of task difficulty.
Proposition 1.
Consider a batch of tasks $\mathcal{B}:=\{\mathcal{T}^{k}\}_{k=1}^{K}$ with ground-truth success probabilities $\{p(\mathcal{T}^{k}),p_{w}^{\mathrm{ref}}(\mathcal{T}^{k}),p_{s}^{\mathrm{ref}}(\mathcal{T}^{k})\}_{k=1}^{K}$. For any task $\mathcal{T}\notin\mathcal{B}$, define the linear interpolation estimator
$\tilde{p}(\mathcal{T})=\mu(\mathcal{B})\,p_{s}^{\mathrm{ref}}(\mathcal{T})+\bigl(1-\mu(\mathcal{B})\bigr)\,p_{w}^{\mathrm{ref}}(\mathcal{T}),$
where
$\mu(\mathcal{B})=\frac{\frac{1}{K}\sum_{k=1}^{K}p(\mathcal{T}^{k})-\frac{1}{K}\sum_{k=1}^{K}p_{w}^{\mathrm{ref}}(\mathcal{T}^{k})}{\frac{1}{K}\sum_{k=1}^{K}p_{s}^{\mathrm{ref}}(\mathcal{T}^{k})-\frac{1}{K}\sum_{k=1}^{K}p_{w}^{\mathrm{ref}}(\mathcal{T}^{k})}.$
Then $\tilde{p}(\mathcal{T})=p(\mathcal{T})$ for all $\mathcal{T}\notin\mathcal{B}$ if and only if Assumption 1 holds.
Proof.
(If) Assume Assumption 1 holds. Let $\mu^{\star}$ be the common interpolation coefficient for all tasks. Then for every $\mathcal{T}^{k}\in\mathcal{B}$,
$p(\mathcal{T}^{k})=\mu^{\star}p_{s}^{\mathrm{ref}}(\mathcal{T}^{k})+(1-\mu^{\star})p_{w}^{\mathrm{ref}}(\mathcal{T}^{k}).$
Taking averages over $\mathcal{B}$, we obtain
$\frac{1}{K}\!\sum_{k=1}^{K}p(\mathcal{T}^{k})=\mu^{\star}\frac{1}{K}\!\sum_{k=1}^{K}p_{s}^{\mathrm{ref}}(\mathcal{T}^{k})+(1-\mu^{\star})\frac{1}{K}\!\sum_{k=1}^{K}p_{w}^{\mathrm{ref}}(\mathcal{T}^{k}).$
Rearranging gives $\mu(\mathcal{B})=\mu^{\star}$.
Now for any $\mathcal{T}\notin\mathcal{B}$, Assumption 1 implies
$p(\mathcal{T})=\mu^{\star}p_{s}^{\mathrm{ref}}(\mathcal{T})+(1-\mu^{\star})p_{w}^{\mathrm{ref}}(\mathcal{T}).$
Since $\mu(\mathcal{B})=\mu^{\star}$, the estimator satisfies
$\tilde{p}(\mathcal{T})=p(\mathcal{T}).$
Thus linear interpolation is perfect under the assumption.
[…]
$p(\mathcal{T}^{i})=\mu(\mathcal{B})\,p_{s}^{\mathrm{ref}}(\mathcal{T}^{i})+\bigl(1-\mu(\mathcal{B})\bigr)\,p_{w}^{\mathrm{ref}}(\mathcal{T}^{i}),\quad i=1,2.$
Solving for $\mu(\mathcal{B})$ in each equation yields
$\frac{p(\mathcal{T}^{1})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{1})}{p_{s}^{\mathrm{ref}}(\mathcal{T}^{1})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{1})}\;=\;\mu(\mathcal{B})\;=\;\frac{p(\mathcal{T}^{2})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{2})}{p_{s}^{\mathrm{ref}}(\mathcal{T}^{2})-p_{w}^{\mathrm{ref}}(\mathcal{T}^{2})}.$
Thus the interpolation ratio is identical for all tasks, which is precisely Assumption 1. This completes the proof. ∎
However, in practice, different tasks do not progress at the same learning pace, and thus Assumption 1 may not hold. This mismatch naturally introduces estimation error into the linear interpolation–based difficulty estimator. Nevertheless, our empirical results show that even with imperfect estimates, the interpolation-based implicit evidence is sufficiently informative to substantially improve training efficiency within BOTS. A more fine-grained theoretical analysis of interpolation error under richer and more realistic assumptions is an interesting direction for future work.
@@ F.3 Interpolation-Based Implicit Evidence
Unlike kernel-based estimators (Sun et al., 2025) (see Appendix F.2 for a brief introduction), which are confined to the convex hull of observed tasks, our interpolation-based estimator naturally supports extrapolation. For example, if the reference pair consists of a strong and a stronger model, the capability coefficient $\mu_{t}(\mathcal{B}_{t})$ defined in Section 3.3 may fall outside $[0,1]$, yielding extrapolated predictions. The clipping step in Equation (5) ensures these estimates remain within the feasible range $[0,1]$.
Additional Practical Benefits of Evaluating Reference Models
We highlight two additional practical benefits of evaluating reference models: (1) as demonstrated in GURU, reference-model evaluations can directly support offline difficulty-aware filtering, removing tasks that are too trivial or too hard before training even begins; and (2) these evaluations can also assist RL algorithms that rely on online difficulty estimation, such as SPO (Xu and Ding, 2025).
@@
The simplicity of our interpolation-based approach, while a key strength, also entails two potential limitations. (i) Expressive power: The linear interpolation assumes a linear relationship between a model’s global capability and its per-task success rate. While our empirical results suggest this is a powerful approximation (see Appendix G.2 for a validation), the true learning dynamics of LLMs may be more complex. (ii) Distributional shift in capability estimation: The capability coefficient $\mu_{t}$ is estimated on the selected batch $\mathcal{B}_{t}$, which is not a uniform sample from the task pool. This introduces a bias, as the model is likely to perform better on this adaptively chosen batch than on the entire dataset. Consequently, $\mu_{t}$ might be an overestimate of the model’s true global capability. However, we argue this bias is not fatal: the primary role of $\mu_{t}$ is to track the progression of the model’s capability, and even a biased estimate can provide a valuable monotonic signal for this purpose, see Section 4.4 for empirical validation. (iii) Potential Failure Mode: If a task exhibits identical success probabilities under both reference models, then the interpolation collapses to that shared value, making the estimated difficulty entirely independent of the current model’s capability.
F.4 Discussion on Future Works
F.4.1 Generalization to Other Reward Distributions
@@
F.4.2 Self-Adaptive Update Rules
@@
F.4.3 Alternative Plug-In for Implicit Evidence
Our interpolation-based estimator provides an extremely lightweight way to generate implicit evidence without additional rollouts, but it is not the only possible choice. In fact, it also comes with an inherent limitation, see discussions in Appendix F.3. More expressive alternatives could be explored, such as kernel-based predictors (Sun et al., 2025), task-embedding regressors, or small auxiliary models trained jointly with the main model. These alternatives may improve predictive accuracy, especially when the reference models poorly bracket the training model’s capability. However, they also introduce a trade-off: richer implicit evidence often requires higher computational and storage costs. Systematically characterizing this trade-off—between accuracy and efficiency in implicit evidence—remains an open research question and a promising direction for future work.
Appendix G Additional Empirical Results
[…]
G.1 Computational Overhead